# Scalable Cosmic AI Inference using Cloud Serverless Computing with FMI

### v2

@@ Abstract
Large-scale astronomical image data processing and prediction is essential for astronomers, providing crucial insights into celestial objects, the universe’s history, and its evolution. While modern deep learning models offer high predictive accuracy, they often demand substantial computational resources, making them resource-intensive and limiting accessibility. We introduce the Cloud-based Astronomy Inference (CAI) framework to address these challenges. This scalable solution integrates pre-trained foundation models with serverless cloud infrastructure through a Function-as-a-Service (FaaS) Message Interface (FMI). CAI enables efficient and scalable inference on astronomical images without extensive hardware. Using a foundation model for redshift prediction as a case study, our extensive experiments cover user devices, HPC (High-Performance Computing) servers, and Cloud. CAI’s significant scalability improvement on large data sizes provides an accessible and effective tool for the astronomy community. The code is accessible at https://github.com/UVA-MLSys/AI-for-Astronomy.
@@ 1 Introduction
Astronomical images are vital to modern astrophysics, offering key insights into celestial objects, such as their shape Willett et al. (2013), distance Hubble (1929), and other fundamental characteristics that define our understanding of the universe. Large surveys like the Dark Energy Spectroscopic Instrument (DESI) Fan et al. (2019) and the Sloan Digital Sky Survey (SDSS) Gunn et al. (1998) provide extensive image datasets with unique attributes. For example, SDSS images consist of five spectral bands—u, g, r, i, and z—each focused on specific wavelengths: Ultraviolet (u) at 3543 Å, Green (g) at 4770 Å, Red (r) at 6231 Å, Near Infrared (i) at 7625 Å, and Infrared (z) at 9134 Å Murrugarra-Llerena and Hirata (2017).
Deep learning foundation models trained on large-scale astronomical image datasets have proven to be powerful tools for improving the accuracy and efficiency of tasks such as redshift prediction, morphology classification, and similarity search Hayat et al. (2021); Lanusse et al. (2023); Fathkouhi and Fox (2024). Figure 5 highlights several notable deep-learning architectures that leverage astronomical images for prediction. For example, AstroCLIP Lanusse et al. (2023) uses an image transformer with a 307M-parameter encoder, pretrained on approximately 197K images. AstroMAE Fathkouhi and Fox (2024) combines frozen pretrained weights with fine-tunable parameters, totaling 10.4M learnable parameters. Similarly, Henghes et al. (2022) trained their model on 1 million SDSS images using 7.8M parameters. In comparison, AstroMAE, trained on about 650K images, showed significant performance improvements over the Henghes et al. (2022) model in redshift prediction.
While these models yield impressive results, their large parameter counts require substantial computational resources, making both training and inference resource-intensive. The memory and computing resources on a standalone device become a limitation for users to run on large datasets and foundation models. There is a need for advanced infrastructure to overcome the limit of high-performance inference accessibility for many users Neely (2021). Serverless computing is a recently popular Function-as-a-Service (FaaS) Li et al. (2022), that lets developers write cloud functions in high-level languages (e.g. Python) and takes care of the complex infrastructure management itself.
This study proposes a highly scalable serverless computing framework for using AWS AWS (2024) to enhance the accessibility of a pretrained foundation model for astronomical images. This effectively reduces the computational demands on individual users. In summary, we offer the following contributions:
@@
Figure 1: The architecture of masked autoencoder of AstroMAE Fathkouhi and Fox (2024).
@@ 2 Related Work
The immense size of astronomy datasets has made cloud services essential for efficiently storing and processing this data. Faaique (2024) highlighted the significant challenges such massive volumes pose, emphasizing cloud computing as a key solution to managing these issues. Kim and Hahm (2011) investigated requirements for cloud computing used by scientists dealing with the increasing size of big science datasets. Gill and Buyya (2020) accepted the invaluable nature of cloud computing for large astronomy datasets and used an astronomy case study to test their proposed cloud computing model that improves failure management. Khlamov et al. (2022) argued that cloud services are nearly indispensable for managing astronomy datasets, especially those dense in metadata and images. In their study on variable star photometry, they implemented a Software-as-a-Service (SaaS) model to effectively address the substantial storage demands.
De Prado et al. (2014) leveraged cloud computing for the Mosaic tool, enabling the rapid and efficient creation of sky mosaics from images captured across different sky regions. Zhang et al. (2020) implemented the distributed astronomy image processing toolkit called Kira by running Apache Spark on an Amazon EC2 cloud. Their speed-up results and throughput indicate that this parallelized computing method is compatible with astronomy applications. Sen et al. (2022) underscored the minimal scalability concerns of cloud services, noting the effectiveness of cloud-based distributed frameworks for tasks like redshift prediction. Furthermore, Parra-Royón et al. (2024) leveraged Function-as-a-Service (FaaS) models and decision-making systems to manage the computationally intensive processing of radio astronomy data.
These studies highlight the essential role of cloud services in overcoming the computational and storage challenges inherent in modern astronomical research. To our knowledge, no prior work has leveraged serverless computing AWS (2024) to enhance the scalability of foundation models for astronomical images.
@@ 3 Problem Statement
In astronomical imaging, much of the research has focused on enhancing deep learning model performance, with comparatively less attention to scaling inference capabilities. Although recent foundation models, trained on extensive astronomical image datasets, demonstrate versatility across various tasks, their high parameter count limits usability and scalability due to infrastructure constraints. To address this, a scalable framework is essential for efficient inference on large image volumes without added financial burdens. We introduce Cloud-based Astronomy Inference (CAI), which employs the Function-as-a-Service (FaaS) Message Interface (FMI) Böhringer (2022) to enhance the scalability of foundation models trained on astronomical images. Based on our review, CAI is the first framework specifically designed to address the inference scalability of foundation models in astronomical imaging.
Although dark energy constitutes approximately 95% of the universe’s energy, our understanding of this mysterious force remains profoundly limited. Investigating and exploring its nature requires a large-scale collection of galaxy images, supported by advanced cosmological methods and theories Jones et al. (2024). A cornerstone of these methods is the precise determination of a critical parameter: redshift. Redshift measures how much the light from a celestial object has been stretched, providing crucial insights into the distances of these objects and the expansion of the universe Hubble (1929).
In this study, we assess CAI’s scalability by applying it to redshift prediction. For this purpose, we selected AstroMAE due to its superior performance and because it is pretrained on a larger dataset, which may enhance generalization compared to AstroCLIP. Future work will explore the scalability of additional models discussed in this context.
@@ 4.1 AstroMAE
AstroMAE Fathkouhi and Fox (2024) is a recent foundation model that captures general patterns in galaxy images for redshift prediction. It has two major phases:
@@ 4.1.1 Pretraining:
Figure 1 illustrates the pretraining process of AstroMAE’s masked autoencoder Devlin (2018). The masked autoencoder aims to reconstruct masked patches using unmasked ones. We mask 75% of the embedded patches, the remaining 25% are fed into the encoder. Initially, images are segmented into uniform patches of size 5×8×8 and embedded into 192-dimensional vectors with positional embedding.
The reconstructed masked patches are compared to their original patches, enabling the model to learn meaningful representations. To promote learning of data patterns instead of memorizing patch positions, the embeddings are randomly shuffled before being input to the encoder. Compared to other pretraining methods like contrastive learning Chen et al. (2020), the masked autoencoder does not rely on specific augmentations, which can potentially increase the dataset size and computational demands. In contrast, AstroMAE’s masked autoencoder processes 25% patches, making it significantly more efficient. A crucial step for working on large astronomy data. AstroMAE also uses a modified ViT Wang et al. (2022), that contains a parallel convolutional module and performs even better.
@@ 4.1.2 Fine-tuning:
Table 1: Redshift Prediction Using Various Architectures with AstroMAE Fathkouhi and Fox (2024).
@@
Table 4.1.2 shows the performance of various architectures combining a pretrained encoder, magnitude block, and the inception model. Rows labeled “from-scratch” denote models where both plain-ViT Dosovitskiy (2020) and pcm-ViT Wang et al. (2022) are initialized and trained entirely from scratch. In contrast, the fine-tuning section includes models where pcm-ViT and plain-ViT are pretrained and frozen during the fine-tuning. All modules in Table 4.1.2 adhere to the architecture shown in Figure 2.
The AstroMAE encoder is pretrained on 80% of the data, 10% for validation, and the remaining 10% for fine-tuning. For fine-tuning, this 10% is further split into 70% for training, 10% for validation, and 20% for testing. The results in Table 4.1.2 are based on inference over the 20% testing samples from the fine-tuning model. The results underscore the rationale behind AstroMAE’s architectural choices.
@@ 4.2 Proposed Framework: CAI
Analyzing large astronomical data requires advanced distributed systems to handle concurrent jobs at scale. We propose a novel cloud architecture called Cloud-based Astronomy Inference (CAI), using AWS serverless AWS (2024) techniques to solve this challenge with significant speed-up improvement. Figure 3 shows an overview of the proposed framework. It has the following steps:
Figure 3: CAI framework overview using AWS serverless computing. It uses the AWS S3 bucket for data, code, and result storage. The state machine defines the workflow steps using AWS lambda functions and distributed maps. Parallel execution is achieved through data partitions for almost linear high-performance inference scaling.
@@
    Data Partitioning: The total data is split into smaller sizes so that each lambda function can work with one small partition at a time during distributed processing. We use a maximum of 25MB of data per partition (arbitrarily chosen. Section 5.7 shows results for different partition sizes). This keeps the data loading time low and allows batch data processing. This is important because too large data can potentially get out of memory (AWS Lambda has a maximum of 10GB memory).
@@
    Distributed Processing: This step is where concurrent job processing is done based on each input item from the initialization. It uses the following AWS components:
@@
    • Serverless Container: The container is a custom AWS Lambda runtime created that includes all required software dependencies. The AWS Lambda function performing model inference is executed inside this runtime. Within this container, we also use the FMI AllReduce collective operation to capture the summation of average running time within the executing pairs of AWS Lambda serverless containers. This algorithm is highly optimized and uses a recursive doubling algorithm implemented by the library. Collectives allows for processing time delineation compared to the overall execution time reported by AWS Lambda. Additionally, with collectives, we can analyze the impact of overhead on the overall execution of our Lambda function outside the running Python code.
@@
Figure 4: CAI: The Architecture of Serverless Cloud with FMI for Cosmic AI Inference.
As part of this work, we introduce and include a novel integration with FMI, FaaS Message Interface Böhringer (2022) for communication over Amazon AWS Lambda and collective operations, as illustrated in 4. This is not supported by default in AWS but offers performance benefits related to message passing between the Lambda functions. FMI supports the following collective operations: 1) send/receive, 2) broadcast, 3) gather, 4) scatter, 5) reduce, 6) all reduce, and 7) scan. Additional operations and languages can be easily added based on the abstract design of the library. FMI uses a rendezvous server during the initialization phase to exchange IP addresses and establish point-to-point connections between a pair of AWS Lambda functions. Functions can pass useful messages or perform collective operations during the later stages (e.g., model inference) once direct communication is established.
@@ 5.1 Dataset
The dataset used in this study is prepared following Fathkouhi and Fox (2024) and originates from Pasquet et al. (2019). It has 659,857 galaxy images derived from SDSS DR8 Aihara et al. (2011). Each image has dimensions of 64$\times$64$\times$5 and is annotated with 64 physical properties, including class, metallicity, and age, along with unique IDs for cross-referencing with other SDSS physical property tables. These images were captured using a 2.5-meter dedicated telescope at Apache Point Observatory in New Mexico and were already background-subtracted and photometrically calibrated. Pasquet et al. (2019) further processed the images by resampling them onto a common grid of overlapping frames and applying the Lánczos-3 resampling kernel Duchon (1979) to enhance image quality and reduce artifacts. The images are center-cropped to a final size of 32$\times$32$\times$5, and the magnitude values for each band, along with the redshift, are obtained using the Astroquery library Ginsburg et al. (2019). The resulting dataset, consisting of images and magnitudes as inputs and redshift as the target, is then prepared for model training.

### v3

@@ Abstract
Large-scale astronomical image data processing and prediction are essential for astronomers, providing crucial insights into celestial objects, the universe’s history, and its evolution. While modern deep learning models offer high predictive accuracy, they often demand substantial computational resources, making them resource-intensive and limiting accessibility. We introduce the Cloud-based Astronomy Inference (CAI) framework to address these challenges. This scalable solution integrates pre-trained foundation models with serverless cloud infrastructure through a Function-as-a-Service (FaaS). CAI enables efficient and scalable inference on astronomical images without extensive hardware. Using a foundation model for redshift prediction as a case study, our extensive experiments cover user devices, HPC (High-Performance Computing) servers, and Cloud. Using redshift prediction with the AstroMAE model demonstrated CAI’s scalability and efficiency, achieving inference on a 12.6 GB dataset in only 28 seconds compared to 140.8 seconds on HPC GPUs and 1793 seconds on HPC CPUs. CAI also achieved significantly higher throughput, reaching 18.04 billion bits per second (bps), and maintained near-constant inference times as data sizes increased, all at minimal computational cost (under \$5 per experiment). We also process large-scale data up to 1 TB to show CAI’s effectiveness at scale. CAI thus provides a highly scalable, accessible, and cost-effective inference solution for the astronomy community. The code is accessible at https://github.com/UVA-MLSys/AI-for-Astronomy.
@@ 1 Introduction
Astronomical images are vital to modern astrophysics, offering key insights into celestial objects, such as their shape (Willett et al., 2013), distance (Hubble, 1929), and other fundamental characteristics that define our understanding of the universe. Large surveys like the Dark Energy Spectroscopic Instrument (DESI) (Fan et al., 2019) and the Sloan Digital Sky Survey (SDSS) (Gunn et al., 1998) provide extensive image datasets with unique attributes. For example, SDSS images consist of five spectral bands—u, g, r, i, and z—each focused on specific wavelengths: Ultraviolet (u) at 3543 Å, Green (g) at 4770 Å, Red (r) at 6231 Å, Near Infrared (i) at 7625 Å, and Infrared (z) at 9134 Å (Murrugarra-Llerena and Hirata, 2017).
Deep learning foundation models trained on large-scale astronomical image datasets have proven to be powerful tools for improving the accuracy and efficiency of tasks such as redshift prediction, morphology classification, and similarity search (Hayat et al., 2021; Lanusse et al., 2023; Fathkouhi and Fox, 2024). Figure 5 highlights several notable deep-learning architectures that leverage astronomical images for prediction. For example, AstroCLIP (Lanusse et al., 2023) uses an image transformer with a 307M-parameter encoder, pretrained on approximately 197K images. AstroMAE (Fathkouhi and Fox, 2024) combines frozen pretrained weights with fine-tunable parameters, totaling 10.4M learnable parameters. Similarly, (Henghes et al., 2022) trained their model on 1 million SDSS images using 7.8M parameters. In comparison, AstroMAE, trained on about 650K images, showed significant performance improvements over the (Henghes et al., 2022) model in redshift prediction.
While these models yield impressive results, their large parameter counts require substantial computational resources, making both training and inference resource-intensive. The memory and computing resources on a standalone device become a limitation for users to run on large datasets and foundation models. There is a need for advanced infrastructure to overcome the limit of high-performance inference accessibility for many users (Neely, 2021). Serverless computing is a recently popular Function-as-a-Service (FaaS) (Li et al., 2022), that lets developers write cloud functions in high-level languages (e.g. Python) and takes care of the complex infrastructure management itself.
This study proposes a highly scalable serverless computing framework for using AWS (AWS, 2024) to enhance the accessibility of a pretrained foundation model for astronomical images. This effectively reduces the computational demands on individual users. In summary, we offer the following contributions:
@@
Figure 1: The architecture of masked autoencoder of AstroMAE (Fathkouhi and Fox, 2024).
@@ 2 Related Work
The immense size of astronomy datasets has made cloud services essential for efficiently storing and processing this data. (Faaique, 2024) highlighted the significant challenges such massive volumes pose, emphasizing cloud computing as a key solution to managing these issues. (Kim and Hahm, 2011) investigated requirements for cloud computing used by scientists dealing with the increasing size of big science datasets. (Gill and Buyya, 2020) accepted the invaluable nature of cloud computing for large astronomy datasets and used an astronomy case study to test their proposed cloud computing model that improves failure management. (Khlamov et al., 2022) argued that cloud services are nearly indispensable for managing astronomy datasets, especially those dense in metadata and images. In their study on variable star photometry, they implemented a Software-as-a-Service (SaaS) model to effectively address the substantial storage demands.
(De Prado et al., 2014) leveraged cloud computing for the Mosaic tool, enabling the rapid and efficient creation of sky mosaics from images captured across different sky regions. (Zhang et al., 2020) implemented the distributed astronomy image processing toolkit called Kira by running Apache Spark on an Amazon EC2 cloud. Their speed-up results and throughput indicate that this parallelized computing method is compatible with astronomy applications. (Sen et al., 2022) underscored the minimal scalability concerns of cloud services, noting the effectiveness of cloud-based distributed frameworks for tasks like redshift prediction. Furthermore, (Parra-Royón et al., 2024) leveraged Function-as-a-Service (FaaS) models and decision-making systems to manage the computationally intensive processing of radio astronomy data.
These studies highlight the essential role of cloud services in overcoming the computational and storage challenges inherent in modern astronomical research. To our knowledge, no prior work has leveraged serverless computing (AWS, 2024) to enhance the scalability of foundation models for astronomical images.
@@ 3 Problem Statement
In astronomical imaging, much of the research has focused on enhancing deep learning model performance, with comparatively less attention to scaling inference capabilities. Although recent foundation models, trained on extensive astronomical image datasets, demonstrate versatility across various tasks, their high parameter count limits usability and scalability due to infrastructure constraints. To address this, a scalable framework is essential for efficient inference on large image volumes without added financial burdens. We introduce Cloud-based Astronomy Inference (CAI), which employs the Function-as-a-Service (FaaS) (Böhringer, 2022) to enhance the scalability of foundation models trained on astronomical images. Based on our review, CAI is the first framework specifically designed to address the inference scalability of foundation models in astronomical imaging.
Although dark energy constitutes approximately 95% of the universe’s energy, our understanding of this mysterious force remains profoundly limited. Investigating and exploring its nature requires a large-scale collection of galaxy images, supported by advanced cosmological methods and theories (Jones et al., 2024). A cornerstone of these methods is the precise determination of a critical parameter: redshift. Redshift measures how much the light from a celestial object has been stretched, providing crucial insights into the distances of these objects and the expansion of the universe (Hubble, 1929).
In astronomical data analysis, the choice of computing infrastructure has a significant impact on efficiency, scalability, cost, and accessibility. Personal computing devices, such as laptops and desktops, provide the most straightforward and accessible environment for initial exploration and analysis of limited scale. These devices have low upfront costs and minimal setup requirements, making them suitable for preliminary testing, educational purposes, or small datasets. However, limited memory and computational capabilities severely restrict their use for larger datasets or complex models. In our experiments, personal devices were unable to handle datasets larger than 8GB, underscoring their limited suitability for extensive astronomical data analysis. High-performance computing (HPC) CPUs offer robust computational power and ample memory, enabling researchers to process large datasets effectively. These systems are reliable, capable of extensive batch processing, and suited to complex computational tasks that may not easily parallelize on GPUs. GPUs within HPC environments significantly accelerate deep learning tasks due to their parallel processing architecture, which excels at inference workloads involving matrix operations typical of neural networks. HPC GPUs achieved much faster inference (140.8 s) compared to CPUs. GPUs are ideal for intensive model training and moderately large-scale inference tasks. However, GPU-based HPC systems come with high fixed and variable costs, as well as resource allocation complexities, making them less accessible to individual researchers or smaller institutions.
In this study, we evaluate the CAI scalability by applying it to the prediction of redshifts. For this purpose, we selected AstroMAE due to its superior performance and because it is pre-trained on a larger dataset. Our proposed CAI framework leverages serverless cloud computing infrastructure to provide an optimal balance of scalability, cost-efficiency, and accessibility. CAI’s linear scalability and minimal computational cost offer substantial advantages over traditional infrastructure. CAI eliminates upfront costs, complex maintenance, and resource management tasks associated with HPC clusters.
Future work will explore the scalability of additional models discussed in this context.
@@ 4.1 AstroMAE
AstroMAE (Fathkouhi and Fox, 2024) is a recent foundation model that captures general patterns in galaxy images for redshift prediction. It has two major phases:
@@ 4.1.1 Pretraining:
Figure 1 illustrates the pretraining process of AstroMAE’s masked autoencoder (Devlin, 2018). The masked autoencoder aims to reconstruct masked patches using unmasked ones. We mask 75% of the embedded patches, the remaining 25% are fed into the encoder. Initially, images are segmented into uniform patches of size 5×8×8 and embedded into 192-dimensional vectors with positional embedding.
The reconstructed masked patches are compared to their original patches, enabling the model to learn meaningful representations. To promote learning of data patterns instead of memorizing patch positions, the embeddings are randomly shuffled before being input to the encoder. Compared to other pretraining methods like contrastive learning (Chen et al., 2020), the masked autoencoder does not rely on specific augmentations, which can potentially increase the dataset size and computational demands. In contrast, AstroMAE’s masked autoencoder processes 25% patches, making it significantly more efficient. A crucial step for working on large astronomy data. AstroMAE also uses a modified ViT (Wang et al., 2022), that contains a parallel convolutional module and performs even better.
@@ 4.1.2 Fine-tuning:
Table 1: Redshift Prediction Using Various Architectures with AstroMAE (Fathkouhi and Fox, 2024).
@@
Table 1 shows the performance of various architectures combining a pretrained encoder, magnitude block, and the inception model. Rows labeled “from-scratch” denote models where both plain-ViT (Dosovitskiy, 2020) and pcm-ViT (Wang et al., 2022) are initialized and trained entirely from scratch. In contrast, the fine-tuning section includes models where pcm-ViT and plain-ViT are pretrained and frozen during the fine-tuning. All modules in Table 1 adhere to the architecture shown in Figure 2.
The AstroMAE encoder is pretrained on 80% of the data, 10% for validation, and the remaining 10% for fine-tuning. For fine-tuning, this 10% is further split into 70% for training, 10% for validation, and 20% for testing. The results in Table 1 are based on inference over the 20% testing samples from the fine-tuning model. The results underscore the rationale behind AstroMAE’s architectural choices.
@@ 4.2 Proposed Framework: CAI
Analyzing large astronomical data requires advanced distributed systems to handle concurrent jobs at scale. We propose a novel cloud architecture called Cloud-based Astronomy Inference (CAI), using AWS Lambda functions (AWS, 2024) to solve this challenge with significant speed-up improvement. Figure 3 shows an overview of the proposed framework. It has the following steps:
Figure 3: CAI framework overview using AWS Lambda Functions. It uses an AWS S3 bucket for data, code, and result storage. The state machine defines the workflow execution steps using AWS Lambda functions and distributed maps. Parallel execution is achieved through data partitions for almost linear high-performance inference scaling.
@@
    Data Partitioning: The total data is split into smaller sizes so that each lambda function can work with one small partition at a time during distributed processing. We use a maximum of 25MB of data per partition (arbitrarily chosen. Section 5.7 shows results for different partition sizes. This keeps the data loading time to a minimum and allows batch data processing. This is important because too large data can potentially run out of memory (AWS Lambda has a maximum of 10,240 MB memory). To correctly size the experiments, we empirically executed a series of experiments to determine the optimal image data size for processing, based on the maximum lambda size. Based on this analysis, Figure 4 illustrates the dataset and model memory usage, considering the partition data size in MB. However, this approach has a limitation compared to other serverless solutions like AWS Fargate (Serverless Docker on AWS ECS) or Apache OpenWhisk. The upper bound of available memory is a constraint in this approach. In contrast, AWS Lambda functions impose upper bounds on CPU and memory availability. As mentioned earlier, the memory limit for AWS Lambda functions is 10,240 MB. The CPU scaling and resource allocation of an AWS Lambda function are directly proportional to its memory configuration.
@@
    Distributed Processing: This step is where concurrent job processing is performed based on each input item from the initialization. It uses the following AWS components:
@@
    • Serverless Container: The container is a custom AWS Lambda runtime created that includes all required software dependencies. The AWS Lambda function performing model inference is executed inside this runtime.
@@
Figure 4: AWS Lambda Memory Usage by Partition Data Size. We empirically size the dataset based on the partition data size in MB.
@@ 5.1 Dataset
The dataset used in this study is prepared following Fathkouhi and Fox (2024) and originates from Pasquet et al. (2019). It has 659,857 galaxy images derived from SDSS DR8 (Aihara et al., 2011). Each image has dimensions of $\times64$$\times$5 and is annotated with 64 physical properties, including class, metallicity, and age, along with unique IDs for cross-referencing with other SDSS physical property tables. These images were captured using a 2.5-meter dedicated telescope at Apache Point Observatory in New Mexico and were already background-subtracted and photometrically calibrated. Pasquet et al. (2019) further processed the images by resampling them onto a common grid of overlapping frames and applying the Lánczos-3 resampling kernel (Duchon, 1979) to enhance image quality and reduce artifacts. The images are center-cropped to a final size of $\times32$$\times$5, and the magnitude values for each band, along with the redshift, are obtained using the Astroquery library (Ginsburg et al., 2019). The resulting dataset, consisting of images and magnitudes as inputs and redshift as the target, is then prepared for model training.